Circular No. 6985

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1810.8–2609 = RX J1810.7–2609 J. Greiner, Astrophysical Institute, Potsdam; A. J. Castro-Tirado, In-

stituto de Astrofísica de Andalucía, Granada; and T. Boller, Max Planck Insitute for Extraterrestrial Physics, Garching, report: “A 1150-s-exposure ROSAT HRI observation was performed on Mar. 24 of the error box of the transient x-ray burster SAX J1810.8–2609 (IAUC 6838). One strong x-ray source was detected at a mean countrate of 0.5 counts/s, corresponding to roughly 1.5 mCrab level in the band 0.1–2.4 keV (absorbed). The bestfit x-ray position is α = 18^{h}10^{m}44^{s}.5, δ = −26^{o}09′01′′ (equinox 2000.0), with an error radius of ±10′′. The x-ray intensity does not vary by more than a factor of 3, and no x-ray burst is detected. The above position was serendipituously covered by a ROSAT PSPC pointing on 1993 Sept. 10, when SAX J1810.8–2609 was not detected (3σ upper limit of 0.027 PSPC counts/s, corresponding roughly to 0.01 HRI counts/s). It was also not detected during the ROSAT all-sky survey in Sept. 1990. This proves the transient nature of our detected ROSAT HRI source. Assuming a power law of photon index 2.2 and full galactic absorption of 3.7 × 10^{21} cm^{−2}, the unabsorbed flux is 1.0 × 10^{−10} erg cm^{−2} s^{−1} in the band 0.1–2.4 keV (and 3.5 × 10^{−11} erg cm^{−2} s^{−1} in the band 2–10 keV) — i.e., a factor of 4 lower than during the BeppoSAX observations on Mar. 10–12 (IAUC 6843). The HRI position of SAX J1810.8–2609 agrees with the BeppoSAX WFC position (with 3′ error radius) to within 30′′ (IAUC 6838), but it is more than 3′ off the BeppoSAX NFI position (with 1′ error radius; IAUC 6843). Thorough analysis of the ROSAT attitude solution does not reveal any anomaly (courtesy W. Grimm, German Space Operation Center, Oberpfaffenhofen) for our ROSAT HRI observation. Images in the Bessel B and the Gunn r and z bands taken on 1998 Mar. 13 with the Danish 1.5-m telescope at La Silla show several optical objects within the HRI error circle. A comparison image in the z band taken on July 25 by G. Israel (Osservatorio di Roma) and S. Covino (Osservatorio di Brera) does not reveal any variability of these objects down to z = 20. Images are posted

at http://www.obs.aip.de/~jcg/xrt/sax1810m2609.html.”

COMET C/1998 K2 (LINEAR) Visual m_{1} estimates: June 20.25 UT, 12.1 (R. Keen, Mt. Thorodin,

CO, 0.32-m reflector); 28.25, 12.4 (A. Hale, Cloudcroft, NM, 0.41-m reflector); July 1.54, 12.1 (D. A. J. Seargent, Cowra, N.S.W., 0.25-m reflector); 21.45, 12.7 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector).

1998 August 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green